OpenStudy (anonymous):

What about this one... Express the set of real numbers graphed on the number line with an inequality. (The numbers graphed are 7 to -12)

OpenStudy (anonymous):

Such a set would be expressed: \[\left\{ x \in \mathbb{R} | -12 \le x \le 7 \right\}\] This set is inclusive on the boundaries, that is, it includes -12 and 7.